OpenStudy (anonymous):

What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0? a. 4x - y = 13 b. 4x + y = 13 c. 4x + y = -13

OpenStudy (ajspeller):

First you need to find the slope of 4x+y-2=0. The slope of the line you have to create will have the same slope. *parallel lines have the same slope Ax+By=C slope = -(A/B) What is A in the given linear equation ? What is B in the given linear equation ?

OpenStudy (daniellelovee):

Wouldn't it be easier to solve if he/she just substituted and then worked PEMDAS backwards? @ajspeller

OpenStudy (ajspeller):

You could but if the answer is the last option in the multiple choice then worked out the problem 3 times instead of just once. But yes, you are correct that is an option.

OpenStudy (daniellelovee):

oh well I would do that just because one could always go back and check and it'll be easier

OpenStudy (anonymous):

I dont get it =/

OpenStudy (ajspeller):

I recommend knowing how to come to the answer because as some point the questions will not be multiple choice. 4x+y=2 Ax+By=2 Slope = -(A/B) = -(4/1)=-4 Then use the point slope formula. (y-y_1)=m(x-x_1) (x_1,y_1) -> (4,-3) (y-(-3))=-4(x-4) (y+3)=-4x+16 y=-4x+13 4x+y=13 --> Solution